OpenStudy (anonymous):

Simplify Each Radical expression. Squareroot45n^5 or V45n5.

Parth (parthkohli):

\[\textbf{HINT}\]\[\sqrt{45n^5} = \sqrt{5 \times 9 \times n^2 \times n^2 \times n} = \sqrt5 \times \sqrt9 \times \sqrt{n^2} \times \sqrt{n^2} \times \sqrt n\]

Parth (parthkohli):

Can you figure the rest out?

OpenStudy (aravindg):

exactly!! take the perfect squares out of the root

OpenStudy (anonymous):

I have no idea how to do them :(

Parth (parthkohli):

...

OpenStudy (aravindg):

can u tell me the perfect squares you knw ?

OpenStudy (anonymous):

do we have to solve the numbers?